OpenStudy (amonoconnor):

I am currently working with Applications of Derivatives, and need clarification on something. The problem: "y = x^2 - x - 1; Determine the following: 1. Inflection Point(s) 2. All Minima & Maxima 3. Intervals where graph is Increasing & Decreasing 4. Intervals of Concave Ups & Downs" So, I took the 2nd derivative of the equation, to determine the Inflection Point and got a single integer- no term containing "x". So what's the I.P.? :)

OpenStudy (anonymous):

it is a quadratic, you don't need calculus for any of this

OpenStudy (amonoconnor):

Sorry for the abbreviations and poor grammar towards the end there- O.S. wouldn't let me type any more characters.

OpenStudy (anonymous):

\[y = x^2 - x - 1\] is a parabolas that opens up

OpenStudy (anonymous):

the minimum value is the second coordinate of the vertex since it is a parabola it is always concave up

OpenStudy (anonymous):

the second derivative, if you really want it, is \(y''=2\) a positive constant that tells you it is always concave up, but you knew that already

OpenStudy (anonymous):

oh also that tells you there is no inflection point, it never changes concavity

OpenStudy (amonoconnor):

So... Like literally the inflection point doesn't exist?

OpenStudy (amonoconnor):

Oh, I get it!!!

OpenStudy (anonymous):

not to belabor the point, but you know all about quadratics without any calculus at all it is decreasing unit you get to the first coordinate of the vertex, which in your case is \(\frac{1}{2}\) then increasing from there on

OpenStudy (anonymous):



OpenStudy (amonoconnor):

Yep, I get the basic structure of parabolas... I guess I just didn't know what to do, on paper, with just 2 since that derivative is normally, well, supposed to be set equal to zero. :) So in this case... is 2 the y-value of the coordinate for the Absolute Minimum?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \(-\frac{-1}{2\times 1}=\frac{1}{2}\)

OpenStudy (anonymous):

second coordinate of the vertex is what you get for \(y\) when you replace \(x\) by \(\frac{1}{2}\)

OpenStudy (amonoconnor):

Ahhh... I wish I had you for a teacher.

OpenStudy (amonoconnor):

Thank you very much!